# algorithmic

### v2

@@ 5.2.1 $\ell_{1}$ regularization
We train the neural model under the regularization function $g(x)=\lambda\left\lVert x\right\rVert_{1}$, with $\lambda=10^{-4}$. We selected the following hyper-parameters: $\rho=10\lambda$, $n_{a}=1$, $Q_{k}=10^{-4}I$, $R_{k}=I$ and $P_{0|0}=100I$. We compare the results to different online optimization alternatives: online ADMM [7] with constant matrix $P=10^{-2}I$ (online-ADMM), EKF-ADMM with time-varying $\rho_{k}=10^{\frac{k}{N}-2}\lambda$ (EKF-ADMMtv), and EKF with $\ell_{1}$-regularization [23] (EKF-$\ell_{1}$). The reason for choosing a time-varying $\rho_{k}$ is that fake measurements are usually not accurate initially, so that it is better to start with a higher value of $\frac{1}{\rho_{k}}$ and then decrease it progressively. In addition, we compare with two offline batch algorithms: NAILM [23] and LBFGS [26], the latter using the Python library jax-sysid. Such batch approaches are considered just for performance comparison. The results obtained at the end of the training phase are averaged over $20$ runs and reported in Table 1.
Table 1: Online learning a static model of (1) with $\ell_{1}$ regularization: mean Loss, Mse, sparsity ratio and execution time (standard deviation) obtained over 20 runs.
@@
EKF-ADMMtv provides the lowest loss. This is also true during the training phase, as shown in Figure 2. Note that all the online algorithms consume the dataset only once (1 epoch), except NAILM and LBFGS that run over 150 and 5000 epochs respectively. The slow execution of online-ADMM is due to solving a non-convex optimization problem at each time step, that we solved using the MATLAB function fminunc (quasi-Newton optimizer). LBFGS provides very sparse solutions, even at the cost of a slightly higher loss function, suggesting that it is particularly suited for sparsification. The online learning performance of EKF-ADMM can be also evaluated by looking at the regret function in Figure 3, where it is also apparent that the proposed algorithm improves the solution quality as more samples are provided.
@@ 5.2.2 Bound constraints
Let us now repeat the training under the bound constraints imposed by the regularization function $g(x)=0$ if $x\in\mathcal{C}$ and $g(x)=+\infty$ otherwise, where $\mathcal{C}=\{x\in\mathbb{R}^{n_{x}}:\ |x_{i}|\leq 0.5\}$. We use the hyper-parameters $\rho=1,n_{a}=5,Q_{k}=10^{-4}I,P_{0|0}=100I$, and $R_{k}=I$. In this example, besides the batch solution obtained by running NAILM, we also run a simple clipping step of the Kalman filter (EKF-CLIP), to compare our proposed approach with a naive solution. Results obtained at the end of the training phase and averaged over $20$ runs are reported in Table 2. EKF-ADMM better enforces constraints that EKF-CLIP. Figure 2 shows the performance of the solution during the training phase. Among the online approaches, considering the final Mse, Cv and execution time, EKF-ADMM provides the best quality solution.
Table 2: Online learning a static model (1) with bound constraints: mean (standard deviation) Mse, constraints violation, and execution time averaged over 20 runs.
@@ 5.3 Online learning of a time-varying model
We test now the ability of EKF-ADMM to adapt the same neural network model, under $\ell_{0}$ regularization, when the data-generating system switches as follows:
$y^{m}_{k}=\begin{cases}\frac{z_{1,k}^{2}-e^{\frac{z_{2,k}}{10}}}{3+|z_{1,k}+z_{2,k}|}+r_{k}\;\;k\leq\frac{N}{3}\\ \frac{z_{1,k}^{2}-e^{\frac{z_{2,k}}{2}}}{3+|z_{1,k}+z_{2,k}|}+r_{k}\;\;\frac{N}{3}<k\leq\frac{2N}{3}\\ \frac{0.3\cdot z_{1,k}^{2}-e^{\frac{z_{2,k}}{2}}}{3+|z_{1,k}+z_{2,k}|}+r_{k}\;\;\frac{2N}{3}<k\\ \end{cases}$ (13)
with $N=1.5\cdot 10^{5}$. We evaluate the regret function $R_{f}(N)=\sum_{k=0}^{N-1}(f_{k}(x_{k})+g(x_{k}))-\min_{z_{1},z_{2},z_{3}}\sum_{i=1}^{3}r_{i}(z_{i})$, with $r_{i}(z_{i})=\sum_{k=(i-1)\frac{N}{3}}^{i\frac{N}{3}}(f_{k}(z_{i})+g(z_{i}))$, where $\{x_{k}\}_{k=0}^{N-1}$ is the sequence generated by Algorithm 1. The regularization term is $g(x)=\lambda\left\lVert w\right\rVert_{0}$, with $\lambda=10^{-4}$, and use the EKF-ADMM hyper-parameters $\rho=10^{3}\cdot\lambda$, $p_{a}=1$, $Q_{k}=10^{-4}I$, and $P_{0|0}=100I$. Since the model is now time-varying, we will also use an EKF implementation with forgetting factor $\alpha=0.9$ [12], which simply amounts of scaling the covariance matrix as $\frac{1}{\alpha}P_{k}$ at each step. The resulting regret function is shown in Figure 5. It is apparent that EKF-ADMM can track changes of the underlying data-generating system.
@@ Appendix A Proof of Theorem 1
Starting from $R_{f}(N)$, since $x_{k+1},\nu_{k+1}$ are the optimal solutions of the first two optimization problems in Algorithm 2 and since $w_{k}=w_{k+1}-\rho(x_{k+1}-\nu_{k+1})$ we have that $\nabla f_{k}(x_{k+1})=-(w_{k+1}-\rho(\nu_{k}-\nu_{k+1}))-\eta(P_{k}^{-1}x_{k+1}-P_{k}^{-1}x_{k})$ and $w_{k+1}\in\partial g(\nu_{k+1})$, where $\partial g$ is the subgradient of $g$. Due to the convexity of $f_{k}(\cdot)$ and $g(\cdot)$,
$f_{k}(x_{k+1})-f_{k}(x^{\star})\leq\nabla f_{k}(x_{k+1})^{T}(x_{k+1}-x^{\star})=-w_{k+1}^{T}(x_{k+1}-\nu^{\star})+\frac{\rho}{2}(\left\lVert\nu^{\star}-\nu_{k}\right\rVert^{2}-\left\lVert\nu^{\star}-\nu_{k+1}\right\rVert^{2}+$ (14)
$\left\lVert x_{k+1}-\nu_{k+1}\right\rVert^{2}-\left\lVert x_{k+1}-\nu_{k}\right\rVert^{2})+\frac{\eta}{2}\left(\left\lVert x^{\star}-x_{k}\right\rVert^{2}_{P_{k}^{-1}}-\left\lVert x^{\star}-x_{k+1}\right\rVert^{2}_{P_{k}^{-1}}-\left\lVert x_{k+1}-x_{k}\right\rVert^{2}_{P_{k}^{-1}}\right)$
and
$g(\nu_{k+1})-g(\nu^{\star})$ $\leq w_{k+1}^{T}(\nu_{k+1}-\nu^{\star})$ (15)
Summing Eqs. (14)-(15) together and noticing that $-w_{k+1}^{T}(x_{k+1}-\nu_{k+1})+\frac{\rho}{2}(x_{k+1}-\nu_{k+1})=\frac{1}{2\rho}(\left\lVert w_{k}\right\rVert^{2}-\left\lVert w_{k+1}\right\rVert^{2})$, we obtain:
@@
$\frac{\rho}{2}(\left\lVert\nu^{\star}-\nu_{k}\right\rVert^{2}-\left\lVert\nu^{\star}-\nu_{k+1}\right\rVert^{2})+\frac{\eta}{2}(\left\lVert x^{\star}-x_{k}\right\rVert^{2}_{P_{k}^{-1}}-\left\lVert x^{\star}-x_{k+1}\right\rVert^{2}_{P_{k}^{-1}}-\left\lVert x_{k+1}-x_{k}\right\rVert^{2}_{P_{k}^{-1}})$
Considering that $f_{k}(x_{k})-f_{k}(x_{k+1})\leq\nabla f_{k}(x_{k})^{T}(x_{k}-x_{k+1})\leq\frac{1}{2\alpha\eta}\left\lVert\nabla f_{k}(x_{k})\right\rVert^{2}+\frac{\alpha\eta}{2}\left\lVert x_{k}-x_{k+1}\right\rVert^{2}$, where the second inequality is due to Fenchel-Young’s inequality, and considering Assumption A1.a of the theorem, we have that:
$f_{k}(x_{k})+g(\nu_{k+1})-(f_{k}(x^{\star})+g(\nu^{\star}))\leq\frac{1}{2\rho}(\left\lVert w_{k}\right\rVert^{2}-\left\lVert w_{k+1}\right\rVert^{2})-\frac{\rho}{2}\left\lVert x_{k+1}-\nu_{k}\right\rVert^{2}+\frac{\rho}{2}(\left\lVert\nu^{\star}-\nu_{k}\right\rVert^{2}$
$\left.-\left\lVert\nu^{\star}-\nu_{k+1}\right\rVert^{2}\right)+\frac{1}{2\alpha\eta}\left\lVert\nabla f_{k}(x_{k})\right\rVert^{2}+\frac{\eta}{2}(\left\lVert x^{\star}-x_{k}\right\rVert_{P_{k}^{-1}}^{2}-\left\lVert x^{\star}-x_{k+1}\right\rVert_{P_{k}^{-1}}^{2}).$
Summing from 0 to $N-1$ and considering Assumption A3 we get
@@
$\leq\frac{1}{2\rho}(\left\lVert w_{0}\right\rVert^{2}-\left\lVert w_{N}\right\rVert^{2})+\frac{\rho}{2}(\left\lVert\nu^{\star}-\nu_{0}\right\rVert^{2}-\left\lVert\nu^{\star}-\nu_{N}\right\rVert^{2})+\frac{1}{2\alpha\eta}\sum_{k=0}^{N-1}\left\lVert\nabla f_{k}(x_{k})\right\rVert^{2}+\frac{\eta}{2}\left\lVert x^{\star}-x_{0}\right\rVert^{2}_{P_{0}^{-1}}$
@@
and, therefore, $R_{f}(N)\leq\frac{\rho}{2}\left\lVert\nu^{\star}\right\rVert^{2}$ $+\frac{1}{2\alpha\eta}\sum_{k=0}^{N-1}\left\lVert\nabla f_{k}(x_{k})\right\rVert^{2}+\frac{\eta}{2}\left\lVert x^{\star}\right\rVert^{2}_{P_{0}^{-1}}+\frac{\eta}{2}\sum_{k=1}^{N-1}\left\lVert x^{\star}-x_{k}\right\rVert^{2}_{(P_{k}^{-1}-P_{k-1}^{-1})}$. Because of Assumption A2, $\frac{1}{2}\sum_{k=1}^{N-1}\left\lVert x^{\star}-x_{k}\right\rVert^{2}_{(P_{k}^{-1}-P_{k-1}^{-1})}=\frac{1}{2}\sum_{k=1}^{k_{n}}\left\lVert x^{\star}-x_{k}\right\rVert^{2}_{(P_{k}^{-1}-P_{k-1}^{-1})}\leq M_{k_{n}}$, and taking into account Assumptions A1.b and A1.c we get $R_{f}(N)\leq\frac{\rho D_{\nu}}{2}+\frac{NG_{f}^{2}}{2\alpha\eta}+\eta(D_{x}^{2}+M_{k_{n}})$. Setting $\eta\stackrel{{\scriptstyle\text{def}}}{{=}}\frac{G_{f}\sqrt{N}}{D_{x}\sqrt{2\alpha}}$ and $\rho\stackrel{{\scriptstyle\text{def}}}{{=}}\sqrt{N}$, we get the sublinear regret bound $R_{f}(N)\leq\frac{\sqrt{N}D_{\nu}}{2}+\frac{G_{f}D_{x}\sqrt{N}}{\sqrt{2\alpha}}+\frac{G_{f}\sqrt{N}(D_{x}^{2}+M_{k_{n}})}{D_{x}\sqrt{2\alpha}}$. Considering now $R_{c}(N)$, we can rearrange (16) and consider Assumption A1.d:
$\left\lVert x_{k+1}-\nu_{k}\right\rVert^{2}\leq\frac{2F}{\rho}+\frac{1}{\rho^{2}}(\left\lVert w_{k}\right\rVert^{2}-\left\lVert w_{k+1}\right\rVert^{2})+(\left\lVert\nu^{\star}-\nu_{k}\right\rVert^{2}-\left\lVert\nu^{\star}-\nu_{k+1}\right\rVert^{2})+\frac{\eta}{\rho}(\left\lVert x^{\star}-x_{k}\right\rVert_{P_{k}^{-1}}^{2}$
$-\left\lVert x^{\star}-x_{k+1}\right\rVert_{P_{k}^{-1}}^{2}-\left\lVert x_{k+1}-x_{k}\right\rVert_{P_{k}^{-1}}^{2}).$
Summing from 0 to $N-1$, we get:

### v3

@@ 5.2.1 $\ell_{1}$ regularization
We train the neural model under the regularization function $g(x)=\lambda\left\lVert x\right\rVert_{1}$, with $\lambda=10^{-4}$. We selected the following hyper-parameters: $\rho=10\lambda$, $n_{a}=1$, $Q_{k}=10^{-4}I$, $R_{k}=I$ and $P_{0|0}=100I$. We compare the results to different online optimization alternatives: online ADMM [13] with constant matrix $P=10^{-2}I$ (online-ADMM), EKF-ADMM with time-varying $\rho_{k}=10^{\frac{k}{N}-2}\lambda$ (EKF-ADMMtv), EKF with $\ell_{1}$-regularization [8] (EKF-$\ell_{1}$) and SMIDAS [20] with learning rate $\eta=5\cdot 10^{-2}$. Notice that $P_{0|0},R_{k}$ and $Q_{k}$ are chosen by manually tuning EKF-$\ell_{1}$ and then used for all other approaches, while all the remaining method-specific hyperparameters have been manually tuned to maximize their performance. The reason for choosing a time-varying $\rho_{k}$ is that fake measurements are usually not accurate initially, so that it is better to start with a higher value of $\frac{1}{\rho_{k}}$ and then decrease it progressively. In addition, we compare with two offline batch algorithms: NAILM [21] and LBFGS [22], the latter using the Python library jax-sysid. Note that all the online algorithms consume the dataset only once (1 epoch), except NAILM and LBFGS that run over 150 and 5000 epochs respectively. The results obtained at the end of the training phase are reported in Table 1.
Table 1: Online learning a static model of (1) with $\ell_{1}$ regularization: mean (standard deviation) Loss, Mse, sparsity ratio and execution time obtained over 20 runs.
@@
SMIDAS [20] $89.93$ ($36.3$) $84.67$ ($36.6$) $61.71$ ($9.98$) $4.11$ ($1.11$)
Among the online approaches, EKF-ADMMtv provides the lowest loss: this is also true during the training phase, as shown in Figure 2. The online learning performance of EKF-ADMM can be also evaluated by looking at the regret function in Figure 3, where it is also apparent that, compared to the other approaches, the proposed algorithm improves the solution quality faster.
@@ 5.2.2 Bound constraints
Let us now repeat the training under the bound constraints imposed by the regularization function $g(x)=0$ if $x\in\mathcal{C}$ and $g(x)=+\infty$ otherwise, where $\mathcal{C}=\{x\in\mathbb{R}^{n_{x}}:\ |x_{i}|\leq 0.5\}$. We use the hyper-parameters $\rho=1,n_{a}=5,Q_{k}=10^{-4}I,P_{0|0}=100I$, and $R_{k}=I$. In this example, we also compare with a simple clipping step of the Kalman filter (EKF-CLIP) and an online projected gradient method (OGD-proj) [4] with learning rate $\eta=10^{-3}$. Matrices $P_{0|0},R_{k}$ and $Q_{k}$ are the same as for the $\ell_{1}$ example and are in common for all the Kalman-like approaches, while all the other hyperparameters have been manually re-tuned to optimize performance in the bounds constraints example. Results obtained at the end of the training phase are reported in Table 2. Among the online approaches, considering the final Mse, Cv and execution time, EKF-ADMM provides the best quality solution, while OGD-proj attains similar performance but at a slower pace. Figure 4 shows the performance of the solution during the training phase.
We have noticed that the choice of $\rho$ is crucial to obtain good adaptation performance: while for some values the convergence is slow and not satisfactory, with limited effort and a proper manual tuning, we were easily able to find values with good performance in all the examples that we have shown here.
Table 2: Online learning a static model (1) with bound constraints: mean (standard deviation) Mse, constraints violation, and execution time obtained over 20 runs.
@@ 5.3 Online learning of a time-varying model
We test now the ability of EKF-ADMM to adapt the same neural network model, under $\ell_{0}$ regularization, when the data-generating system switches as follows: $y^{m}_{k}=\frac{z_{k,1}^{2}-e^{\frac{z_{k,2}}{10}}}{3+|z_{k,1}+z_{k,2}|}+r_{k}$ if $k\leq\frac{N}{3}$, $y^{m}_{k}=\frac{z_{k,1}^{2}-e^{\frac{z_{k,2}}{2}}}{3+|z_{k,1}+z_{k,2}|}+r_{k}$ if $\frac{N}{3}<k\leq\frac{2N}{3}$ and $y^{m}_{k}=\frac{0.3\cdot z_{k,1}^{2}-e^{\frac{z_{k,2}}{2}}}{3+|z_{k,1}+z_{k,2}|}+r_{k}$ if $\frac{2N}{3}<k$, with $N=1.5\cdot 10^{5}$, $z_{k,i}\sim\mathcal{U}_{[-10,10]}$ and $r_{k}\sim{\mathcal{N}}(0,10^{-2})$. We evaluate the regret function $R_{f}(N)=\sum_{k=0}^{N-1}(f_{k}(x_{k})+g(x_{k}))-\min_{z_{1},z_{2},z_{3}}\sum_{i=1}^{3}r_{i}(z_{i})$, with $r_{i}(z_{i})=\sum_{k=(i-1)\frac{N}{3}}^{i\frac{N}{3}}(f_{k}(z_{i})+g(z_{i}))$, where $\{x_{k}\}_{k=0}^{N-1}$ is the sequence generated by Algorithm 1. The regularization term is $g(x)=\lambda\left\lVert w\right\rVert_{0}$, with $\lambda=10^{-4}$, and we use the EKF-ADMM hyper-parameters $\rho=10^{3}\cdot\lambda$, $p_{a}=1$, $Q_{k}=10^{-4}I$, and $P_{0|0}=100I$, which were manually tuned by optimizing the EKF performance. Since the model is now time-varying, we will also use an EKF implementation with forgetting factor $\alpha=0.9$ [9]. The resulting regret function is shown in Figure 5. It is apparent that EKF-ADMM can effectively track changes of the underlying data-generating system.
@@ Appendix A Proof of Theorem 1
Starting from $R_{f}(N)$, since $x_{k+1},\nu_{k+1}$ are the optimal solutions of the first two optimization problems in Algorithm 2 and since $w_{k}=w_{k+1}-\rho(x_{k+1}-\nu_{k+1})$ we have that $\nabla f_{k}(x_{k+1})=-(w_{k+1}-\rho(\nu_{k}-\nu_{k+1}))-\eta(P_{k}^{-1}x_{k+1}-P_{k}^{-1}x_{k})$ and $w_{k+1}\in\partial g(\nu_{k+1})$, where $\partial g$ is the subgradient of $g$. Due to the convexity of $f_{k}(\cdot)$ and $g(\cdot)$, $f_{k}(x_{k+1})-f_{k}(x^{\star})\leq\nabla f_{k}(x_{k+1})^{T}(x_{k+1}-x^{\star})=-w_{k+1}^{T}(x_{k+1}-\nu^{\star})+\frac{\rho}{2}(\left\lVert\nu^{\star}-\nu_{k}\right\rVert^{2}-\left\lVert\nu^{\star}-\nu_{k+1}\right\rVert^{2}+\left\lVert x_{k+1}-\nu_{k+1}\right\rVert^{2}-\left\lVert x_{k+1}-\nu_{k}\right\rVert^{2})+\frac{\eta}{2}(\left\lVert x^{\star}-x_{k}\right\rVert^{2}_{P_{k}^{-1}}-$ $\left\lVert x^{\star}-x_{k+1}\right\rVert^{2}_{P_{k}^{-1}}-\left\lVert x_{k+1}-x_{k}\right\rVert^{2}_{P_{k}^{-1}})$, where the equality derives from a succession of square completions and by recalling that $x^{\star}=\nu^{\star}$, and $g(\nu_{k+1})-g(\nu^{\star})\leq w_{k+1}^{T}(\nu_{k+1}-\nu^{\star})$. Summing the two inequalities together and noticing that $-w_{k+1}^{T}(x_{k+1}-\nu_{k+1})+\frac{\rho}{2}\lVert x_{k+1}-\nu_{k+1}\rVert^{2}=\frac{1}{2\rho}(\left\lVert w_{k}\right\rVert^{2}-\left\lVert w_{k+1}\right\rVert^{2})$, we obtain:
@@
$+\frac{\rho}{2}(\left\lVert\nu^{\star}-\nu_{k}\right\rVert^{2}-\left\lVert\nu^{\star}-\nu_{k+1}\right\rVert^{2})+\frac{\eta}{2}(\left\lVert x^{\star}-x_{k}\right\rVert^{2}_{P_{k}^{-1}}-\left\lVert x^{\star}-x_{k+1}\right\rVert^{2}_{P_{k}^{-1}}-\left\lVert x_{k+1}-x_{k}\right\rVert^{2}_{P_{k}^{-1}})$
Considering that $f_{k}(x_{k})-f_{k}(x_{k+1})\leq\nabla f_{k}(x_{k})^{T}(x_{k}-x_{k+1})\leq\frac{1}{2\alpha\eta}\left\lVert\nabla f_{k}(x_{k})\right\rVert^{2}+\frac{\alpha\eta}{2}\left\lVert x_{k}-x_{k+1}\right\rVert^{2}$, where the second inequality is due to Fenchel-Young’s inequality, and considering Assumption A1.a of the theorem, we have that $f_{k}(x_{k})+g(\nu_{k+1})-(f_{k}(x^{\star})+g(\nu^{\star}))\leq\frac{1}{2\rho}(\left\lVert w_{k}\right\rVert^{2}-\left\lVert w_{k+1}\right\rVert^{2})-\frac{\rho}{2}\left\lVert x_{k+1}-\nu_{k}\right\rVert^{2}+\frac{\rho}{2}(\left\lVert\nu^{\star}-\nu_{k}\right\rVert^{2}-\left\lVert\nu^{\star}-\nu_{k+1}\right\rVert^{2})+\frac{1}{2\alpha\eta}\left\lVert\nabla f_{k}(x_{k})\right\rVert^{2}+\frac{\eta}{2}(\left\lVert x^{\star}-x_{k}\right\rVert_{P_{k}^{-1}}^{2}-\left\lVert x^{\star}-x_{k+1}\right\rVert_{P_{k}^{-1}}^{2})$. Summing from 0 to $N-1$ and considering Assumption A3 we get
@@
$\leq\frac{1}{2\rho}(\left\lVert w_{0}\right\rVert^{2}-\left\lVert w_{N}\right\rVert^{2})+\frac{\rho}{2}(\left\lVert\nu^{\star}-\nu_{0}\right\rVert^{2}-\left\lVert\nu^{\star}-\nu_{N}\right\rVert^{2})+\frac{1}{2\alpha\eta}\sum_{k=0}^{N-1}\left\lVert\nabla f_{k}(x_{k})\right\rVert^{2}+\frac{\eta}{2}\left\lVert x^{\star}-x_{0}\right\rVert^{2}_{P_{0}^{-1}}+$
@@
and, therefore, $R_{f}(N)\leq\frac{\rho}{2}\left\lVert\nu^{\star}\right\rVert^{2}$ $+\frac{1}{2\alpha\eta}\sum_{k=0}^{N-1}\left\lVert\nabla f_{k}(x_{k})\right\rVert^{2}+\frac{\eta}{2}\left\lVert x^{\star}\right\rVert^{2}_{P_{0}^{-1}}+\frac{\eta}{2}\sum_{k=1}^{N-1}\left\lVert x^{\star}-x_{k}\right\rVert^{2}_{(P_{k}^{-1}-P_{k-1}^{-1})}$. Because of Assumption A2, $\frac{1}{2}\sum_{k=1}^{N-1}\left\lVert x^{\star}-x_{k}\right\rVert^{2}_{(P_{k}^{-1}-P_{k-1}^{-1})}=\frac{1}{2}\sum_{k=1}^{k_{n}}\left\lVert x^{\star}-x_{k}\right\rVert^{2}_{(P_{k}^{-1}-P_{k-1}^{-1})}\leq M_{k_{n}}$, and taking into account Assumptions A1.b and A1.c we get $R_{f}(N)\leq\frac{\rho D_{\nu}}{2}+\frac{NG_{f}^{2}}{2\alpha\eta}+\eta(D_{x}^{2}+M_{k_{n}})$. Setting $\eta\stackrel{{\scriptstyle\text{def}}}{{=}}\frac{G_{f}\sqrt{N}}{D_{x}\sqrt{2\alpha}}$ and $\rho\stackrel{{\scriptstyle\text{def}}}{{=}}\sqrt{N}$, we get the sublinear regret bound $R_{f}(N)\leq\frac{\sqrt{N}D_{\nu}}{2}+\frac{G_{f}D_{x}\sqrt{N}}{\sqrt{2\alpha}}+\frac{G_{f}\sqrt{N}(D_{x}^{2}+M_{k_{n}})}{D_{x}\sqrt{2\alpha}}$. Considering now $R_{c}(N)$, we can rearrange (14) and consider Assumption A1.d, $\left\lVert x_{k+1}-\nu_{k}\right\rVert^{2}\leq\frac{2F}{\rho}+\frac{1}{\rho^{2}}(\left\lVert w_{k}\right\rVert^{2}-\left\lVert w_{k+1}\right\rVert^{2})+(\left\lVert\nu^{\star}-\nu_{k}\right\rVert^{2}-\left\lVert\nu^{\star}-\nu_{k+1}\right\rVert^{2})++\frac{\eta}{\rho}\left(\left\lVert x^{\star}-x_{k}\right\rVert_{P_{k}^{-1}}^{2}-\left\lVert x^{\star}-x_{k+1}\right\rVert_{P_{k}^{-1}}^{2}-\left\lVert x_{k+1}-x_{k}\right\rVert_{P_{k}^{-1}}^{2}\right).$ Summing from 0 to $N-1$, we get: